Smart Inc. is going to purchase a $20,000 machine. The vendor offered a term of 2/15 net 45 of payment.

  1. If Smart takes a cash discount when should a company make payment and how much is payment?

  2. If company give up discount when and how much should company pay?

  3. What is the cost of giving up the discount?
Answer #1

2/15 net 45 means If payment is made within 15 days 2% cash dicount is available. In any other case payment shall be done within 45 days.

If the cash discount is availed,

Payment Shall be = 20,000*(100 - 2)% = 20,000 * 98% = $19,600

If the company does not avail the discount,

Payment shall be $20,000

Cost of giving up the discount shall be

Amount foregone = (20,000 - 19,600) = $400

Period = 30 days

(Note : Even though the payment is made within 15 days, he could have $400 profit and the extra amount is paid to delay the payment for 30 days)

Cost of Giving Up = \frac{400}{20,000} * \frac{365}{30} = 24.33 percentage